Enorym:

The graph of a quadratic function passes through the points (a-4, 0) and (a+2, 0), and the coordinates of the vertex are (5, -2). Find the value of a.

sillybilly123:



sillybilly123:

symmetry is your friend

sillybilly123:



Enorym:

How would you do this algebraically?

sillybilly123:

depends upon what you already know @Enorym do you get the original point of the problem?

sillybilly123:

hey Voca !! can you write a write an equation that singularly passes through all of these points?

Vocaloid:

yikes I'll try if a - 4 and a + 2 are zeros, you can write an equation in factored form like y = (x - (a-4))(x - (a+2)) you are given the point (5, -2) so you can plug these in for x and y to find a ^ I hope this works :S

dude:

\(y=c(x−5)2−2\\ because ~it ~passes ~through ~(a-4,0) ~and ~(a+2,0)\\ 0=c(a-4-5)^2-2\\ or ~c(a−9)2=2...(1)\\ ~\\~ and\\ 0=c(a+2−5)2−2\\ ~\\ or c(a−3)2=2...(2)\\ ~\\ dividing (1) by (2)\\ c(a−9)2c(a−3)2=1\\ ~\\ (a−9)2=(a−3)2\\ ~\\ a2−18a+81=a2−6a+9\\ ~\\ a2−18a−a2+6a=9−81\\ ~\\ −12a=−72\\ ~\\ a=6\) Different explanation